OpenStudy (anonymous):

Jane corporation produces fudge bars. Total fixed cost is $55,500. Each package of fudge bars sells for $4.95 with a variable unit cost of $3.10. What is the breakeven point for Jane Corp.

OpenStudy (anonymous):

breakeven means sales = production costs Let the breakeven come at the sale of x units then production cost = 3.10*x + 55,500 = 3.10x + 55,500 sales amount = 4.95*x = 4.95x for breakeven, 4.95x = 3.10x + 55,500 4.95x - 3.10x = 55,500 1.85x = 55,500 x = 55,500/1.85 x = 30,000 the breakeven point is the sale of 30,000 units of fudge bars.